# MirrorCalib: Utilizing Human Pose Information for Mirror-based Virtual Camera Calibration

### v2

@@ Abstract
In this paper, we present the novel task of estimating the extrinsic parameters of a virtual camera relative to a real camera in exercise videos with a mirror. This task poses a significant challenge in scenarios where the views from the real and mirrored cameras have no overlap or share salient features. To address this issue, prior knowledge of a human body and 2D joint locations are utilized to estimate the camera extrinsic parameters when a person is in front of a mirror. We devise a modified eight-point algorithm to obtain an initial estimation from 2D joint locations. The 2D joint locations are then refined subject to human body constraints. Finally, a RANSAC algorithm is employed to remove outliers by comparing their epipolar distances to a predetermined threshold. MirrorCalib is evaluated on both synthetic and real datasets and achieves a rotation error of 0.62°/1.82° and a translation error of 37.33/69.51 mm on the synthetic/real dataset, which outperforms the state-of-art method.
Index Terms:
camera calibration, catadioptric, human pose estimation
I Introduction
@@
Estimating 3D human motion in exercise videos is crucial to many applications including but not limited to the assessment of trainee exercise quality [36], virtual or augmented reality [37], and motion captioning [38, 39]. However, reconstructing 3D human motion from monocular 2D video input is an ill-posed problem, due to depth ambiguity. The inclusion of mirror view in exercise videos (e.g., Figure 1) provides an additional perspective for the instructors. If the geometric relationship between the real view and the virtual view can be established, in principle, it becomes feasible to determine 3D joint positions through triangulation.
@@
Figure 1: Given images of an individual in front of a mirror, after extrinsic parameter calibration of the virtual camera with respect to a real camera, we can reconstruct the 3D human pose.
@@
MirrorCalib has been validated on both synthetic datasets and real datasets. It outperforms the existing methods in the literature which also use 2D human joint locations to calibrate extrinsic parameters in a general multi-view camera setup. Specifically, when intrinsic parameters are known, MirrorCalib achieves a rotation error of 0.62° and a translation error of 37.33 mm on the synthetic dataset and a rotation error of 1.82° and a translation error of 69.51 mm on the real dataset. We have also demonstrated the application of MirrorCalib in 3D human pose estimation with mirrored view and found MirrorCalib contributes to two-third reduction in 3D joint location errors compared to a baseline method that does not perform optimization and outlier removal.
Compared to existing work that calibrates virtual cameras[6, 7, 8, 9], our work differs in two key aspects. First, previous studies[6, 7, 8, 9] often calibrate virtual cameras in catadioptric stereo systems using multiple mirrors. Such systems are not common in daily use. In contrast, MirrorCalib only requires the view of one fixed mirror, which is widely available in gyms and dance studios. Second, since earlier work considers scenarios with more than one mirror, it is easier to obtain overlapping views and rely on visible points to establish correspondences. In contrast, our work utilizes (generally invisible) 2D human joint locations as feature correspondences.
The rest of the paper is organized as follows: a brief review of related work in the literature is provided in Section II, followed by a presentation of the necessary background for our approach in Section III. In Section IV, a detailed explanation of MirrorCalib is provided. Experiment setup, datasets and evaluation results are presented in Section V followed by conclusion in Section VI.
@@
II Related Work
II-A Camera Calibration with Mirrors
Using mirrors to obtain an additional perspective of objects has been exploited in the literature. This approach is applied in estimating camera pose with respect to an object that is not directly visible to the camera and in 3D reconstruction.
In extrinsic camera calibration, incorporating mirrors provides a viable solution when reference objects are outside the camera FOV. For example, Sturm et al. [26] first estimate the virtual camera pose and mirror planes, and then the real camera pose can be obtained by reflecting the virtual view. They find that the planar motion between pairs of virtual views can be regarded as a fixed-axis rotation and the rotation axis can be estimated. All mirror planes should contain this rotation axis and their positions are determined by the rotation axis and the angle between pairs of virtual views. They also demonstrate that a minimum of three virtual views are needed to determine the positions of mirrors and camera poses. Similarly, Rodrigues et al. [1] estimate the pose of a real camera using a minimum of three virtual views and by employing knowledge of fixed-axis rotations. However, their approach differs in computing the mirror direction in a single step without relying on the inference from the axis angle. Without using fixed-axis rotation as constraints, Kumar et al. [3] find out that for each column vector of the rotation matrix of the real camera, the summation of the vector with the corresponding column vector of the rotation matrix of its mirrored camera is orthogonal to the vector connecting a real camera and its mirrored camera. They apply this constraint to solve the real camera pose but require at least five mirrored images. Long et al. [4] also provide a solution that does not rely on the fixed-axis rotation constraint. They formulate the rotation of the real camera and the rotation of its mirrored camera as a rotation averaging problem and solve it using an optimization framework. All the aforementioned methods do not consider the number of known 3D reference points. In case where the number is three, there can be up to four possible solutions for extrinsic parameters. If 3 mirrored images are utilized to recover the real camera pose, 64 = $4\crossproduct 4\crossproduct 4$ combinations of 3D reference points are possible. Takahashi et al. [2] propose an orthogonality constraint to choose the best combination of reference points without explicitly computing the corresponding extrinsic parameters. Instead of planar mirrors, Agrawal and Ramalingam [5] utilize spherical mirrors in their approach to avoid the problem of parallel mirrors configuration. However, planar mirrors are still more common in practice, and our paper only considers planar mirrors.
Catadioptric stereo systems, leveraging multiple mirrors for additional views, have also been applied in 3D reconstruction. This task is close to our scenario. However, the extrinsic parameters of a virtual camera are usually calibrated relative to another virtual camera. Some authors use a single perspective camera and two planar mirrors to build a stereo system [6, 7, 8, 9]. Those works calibrate the virtual cameras using image correspondences like general structure-from-motion methods. After the derivation of epipolar geometry based on mirror constraints, the number of unknown extrinsic calibration parameters can be reduced. In this work, we use 2D human joint correspondences to calculate the extrinsic parameters of the virtual camera with respect to a real camera.
To the best of our knowledge, our paper is the first work that considers calibrating the virtual camera with respect to a real camera in a catadioptric stereo system with only one fixed planar mirror. In contrast, existing methods targeting catadioptric stereo systems either require a mirror to be moved to multiple positions or involve multiple mirrors to obtain overlapping views.
II-B Extrinsic Camera Calibration with Human Pose Information
Prevailing methods of camera calibration utilize chessboard corners or local features like SIFT as corresponding points to recover relative camera poses. When the cameras have little overlapping views (wide baselines), image feature correspondences are difficult to establish. Incorrect feature correspondences can severely degrade camera calibration results. Therefore, methods have been proposed to utilize prior knowledge of humans to calibrate cameras. The studies in [14, 15] use the silhouettes of humans to establish the corresponding points. With the advances in 2D human pose estimation from single images, Puwein et al. [11] propose the first method to use detected 2D human joints in multi-view images as feature points to calculate the extrinsic parameters of the respective cameras. After initial calibration, with the initial 3D joint positions obtained through triangulation, the camera parameters and 3D human poses are further optimized by the application of constraints on smooth motion, ensuring the consistency of detected poses with optical flow, and taking into account the discriminative scores of 2D joints. Similarly, Takahashi et al. [12] present an approach that utilizes a relaxed reprojection error in robust optimization. As most 2D pose estimators produce a confidence map for each joint, the reprojection error is small when the reprojected point is in an area of high confidence. Huang et al. [13] design a VAE-based latent motion prior to jointly optimizing extrinsic parameters and human motion. All of these works use noisy 2D joint estimation to compute initial camera extrinsic parameters and 3D joint positions, and then refine the parameters and human joint positions together through an optimization process.
This work focuses on calibrating virtual camera parameters relative to a real camera, which differs from multi-view camera calibration. Conceptually, methods in [11, 12, 13] can be applied to our problem setting as well. However, the direct application of these methods fails to exploit the geometry constraints imposed by mirrors that can reduce the number of unknowns in the calculation of the fundamental matrix. Moreover, a RANSAC algorithm that leverages epipolar distances to identify the inliers is applied in MirrorCalib.
III Background
@@
In this section, key geometric concepts relevant to estimating virtual camera poses with respect to a real camera are provided and the general normalized eight-point algorithm for fundamental matrix calculation is introduced. It serves as the basis for the proposed eight-point algorithm with mirror constraints in the Methodology section.
III-A Planar Mirror Reflection Geometry
In this paper, we treat the real camera frame as the coordinate system where the optical center of the real camera is the origin. Let the intrinsic parameter matrix of the camera be $\mathbf{K}$. We consider the configuration depicted in Figure 3, in which a perspective camera $C$ is in front of a planar mirror $\pi$ that is uniquely defined by its normal vector $\mathbf{n}$ and a distance $d$ to $C$. Any point $x$ on $C$’s image plane satisfies the condition that $\mathbf{n}^{T}x=d$. The symmetric transformation between a point $p$ and its mirrored point $p^{\prime}$ with respect to $\pi$ is given by:
$\begin{bmatrix}p^{\prime}\\ 1\end{bmatrix}=S\begin{bmatrix}p\\ 1\end{bmatrix},\quad\text{where}\quad S=\begin{bmatrix}\mathbf{I}-2\mathbf{nn}^{T}&2d\mathbf{n}\\ \bm{0}&1\end{bmatrix}$ (7)
Denote by $\tilde{\mathbf{u}}$ and $\tilde{\mathbf{u}}^{\prime}$ the homogeneous coordinates of the point $p$ and the point $p^{\prime}$ projected onto the image planes of the real camera $C$. Since the projection $\tilde{\mathbf{u}}$ of $p$ is given by:
$\alpha\tilde{\mathbf{u}}=\mathbf{K}\begin{bmatrix}\mathbf{I}&\bm{0}\end{bmatrix}\begin{bmatrix}p\\ 1\end{bmatrix}\quad\text{where }\alpha\in\mathbb{R}^{+}\text{is a scale parameter},$ (8)
the projection $\mathbf{u^{\prime}}$ of the mirrored point $p^{\prime}$ is:
$\alpha\tilde{\mathbf{u}}^{\prime}=\mathbf{K}\mathbf{S}\begin{bmatrix}\mathbf{I}&\bm{0}\end{bmatrix}\begin{bmatrix}p\\ 1\end{bmatrix}\text{ }\text{where }\alpha\in\mathbb{R}^{+}\text{is a scale parameter}.$ (9)
@@ III-A Planar Mirror Reflection Geometry
III-B Virtual Camera Projection with Reflective Surfaces and Reflective Epipolar Geometry
The projection $\mathbf{u^{\prime}}$ of the mirrored point $p^{\prime}$ can be seen as the projection $\mathbf{u}^{\prime\prime}$ of the point $p$ on the image plane of a virtual camera $C^{\prime}$. The reference frame of the virtual camera is obtained by reflecting that of the real camera about the mirror $\pi$. It should be noted that the real camera is right-handed i.e., the orientation of the camera’s coordinate system follows the right-hand rule, while the virtual camera is left-handed. As shown in Figure 3, the projection of point $p^{\prime}$ on $C$ is the same as the projection of $p$ onto the image plane of $C’$ ’. Since both the real camera and its virtual counterpart share the same set of intrinsic parameters, the purpose of camera calibration in this context is to estimate the extrinsic parameters between these two camera frames, specifically, the rotation matrix $\mathbf{R}$ and the translation vector $\mathbf{t}$.
From [16], $\tilde{\mathbf{u}}$ and $\tilde{\mathbf{u}}^{\prime}$ satisfy the reflective epipolar constraint:
$(\tilde{\mathbf{u}}^{\prime})^{T}\mathbf{F}\tilde{\mathbf{u}}=0,$ (10)
@@ III-B Virtual Camera Projection with Reflective Surfaces and Reflective Epipolar Geometry
$\mathbf{F}\triangleq\mathbf{K}^{-T}\mathbf{E}\mathbf{K}^{-1}$ (11)
@@
is the reflective essential matrix, where $[\mathbf{n}]_{\times}$ is the skew-symmetric matrix associated with the normal vector $\mathbf{n}$. Note that since the reflective essential matrix is a skew-symmetric matrix, after transformation by the intrinsic camera matrix, the reflective fundamental matrix is also a skew-symmetric matrix.
III-C General Normalized Eight-Point Algorithm
The projections of a point onto the image planes of two cameras related by a fundamental matrix $F$ satisfy Equation (10). For every matched pair $\tilde{\mathbf{u}}=[u,v,1]^{T}$, $\tilde{\mathbf{u}}^{\prime}=[u^{\prime},v^{\prime},1]^{T}$, we have a linear equation of unknown entries in fundamental matrix $\mathbf{F}$. If we denote a row vector $[uu^{\prime},uv^{\prime},u,vu^{\prime},vv^{\prime},v,u^{\prime},v^{\prime},1]^{T}$ by $\mathbf{a}$ and stack all entries of $\mathbf{F}$ in a row-major order to a vector $\mathbf{f}$, Equation (10) is equivalent to $\mathbf{a^{T}f}=0$. Given multiple pairs of ($\tilde{\mathbf{u}}$, $\tilde{\mathbf{u}}^{\prime}$), the system of equations is given by:
$\mathbf{Af}=\mathbf{0},\text{ }\mathbf{||f||}=1.$ (13)
Since the rank of $\mathbf{A}$ is eight [17], a solution can be found with eight matched non-coplanar pairs. When the number of matched pairs exceeds eight, Equation (13) is over-determined. Furthermore, real-world data often contain noise, which may result in infeasibility. Therefore, a least-square optimization can be formulated to minimize the $L_{2}$ norm of $\mathbf{Af}$. The solution to the optimization problem can mitigate the effects of noise and improve the accuracy of the estimation.
Directly applying the classic eight-point algorithm to compute the fundamental matrix is known to be highly sensitive to noise. In [17], it is shown that normalizing the coordinates of the matched points $\tilde{\mathbf{u}}_{i}^{\prime}=[u^{\prime}_{i},v^{\prime}_{i},1]$ and $\tilde{\mathbf{u}}_{i}=[u_{i},v_{i},1]$ can improve noise robustness. Specifically, let
$\hat{\tilde{\mathbf{u}}}^{\prime}=\mathbf{T^{\prime}}\tilde{\mathbf{u}}’$ ’ (14)
$\hat{\tilde{\mathbf{u}}}=\mathbf{T}\tilde{\mathbf{u}}$ (15)
$\mathbf{T}=\begin{bmatrix}s&0&-st_{x}\\ 0&s&-st_{y}\\ 0&0&1\end{bmatrix},\quad\mathbf{T^{\prime}}=\begin{bmatrix}s^{\prime}&0&-s^{\prime}t_{x}^{\prime}\\ 0&s^{\prime}&-st_{y}^{\prime}\\ 0&0&1\end{bmatrix}.$ (16)
Then the original fundamental matrix can be computed by:
$\mathbf{F}=\mathbf{T^{\prime}}^{T}\mathbf{F^{\prime}}\mathbf{T},$ (17)
where $s$ and $s^{\prime}$ are positive scale factors, $(t_{x},t_{y})$ and $(t_{x}^{\prime},t_{y}^{\prime})$ are coordinates of the centroids of all image points of the real view and its virtual view, and $\mathbf{F}^{\prime}$ is the fundamental matrix for normalized image points, respectively. This normalization brings the centroid of the set of points in each image to the origin and normalizes coordinates such that the average distance of points to the origin is $\sqrt{2}$. Then for normalized eight-point algorithm, each row vector of $\mathbf{A}$ in equation (13) is formed by normalized coordinates $[\hat{u}\hat{u}^{\prime},\hat{u}\hat{v}^{\prime},\hat{u},\hat{v}\hat{u}^{\prime},\hat{v}\hat{v}^{\prime},\hat{v},\hat{u}^{\prime},\hat{v}^{\prime},1]^{T}$. And $\mathbf{F}^{\prime}$ can be solved using a least-square optimization algorithm when the number of matched pairs exceeds eight.
IV Methodology
@@
IV-A Eight-Point Algorithm for Solving Reflective Fundamental Matrix
Typically a pair of real cameras have nine unknowns in the fundamental matrix. However, in the case of the reflective fundamental matrix between a real camera and its virtual one, as explained in Section III-B, the fundamental matrix is a skew-symmetric matrix determined by only three unknowns. Therefore, the transformed fundamental matrix after image normalization is:
@@ IV-A Eight-Point Algorithm for Solving Reflective Fundamental Matrix
where $[x_{1},x_{2}...x_{6}]$ are unknowns. The fundamental matrix can be solved through least-square minimization. Specifically, the matrix $\mathbf{\hat{A}}$ is obtained by forming a row vector for each pair of matching points $[-\hat{v}^{\prime}\hat{u}+\hat{u}^{\prime}\hat{v},\hat{u}^{\prime},\hat{v}^{\prime},\hat{u},\hat{v},1]$, where $\hat{u}$’s and $\hat{v}$’s are normalized coordinates. The objective is to minimize:
@@
where $\mathbf{f’}$ ’ is the vector $[x_{1},x_{2}...x_{6}]^{T}$ that determines $\mathbf{F}^{\prime}$. The least-square solution to Equation (19) can be calculated by getting the smallest eigenvalue of $\mathbf{A^{T}A}$. After $\mathbf{F^{\prime}}$ is obtained, we transform it to the original reflective fundamental matrix $\mathbf{F}$ by Equation (17).
IV-B Reflective Essential Matrix Decomposition
@@
Similar to a conventional essential matrix, $\mathbf{R}$ and $\mathbf{t}$ can be obtained from the reflective essential matrix $\mathbf{E}$ using singular value decomposition (SVD). Details of the procedure can be found in the Appendix.
IV-C Estimating Virtual Camera Extrinsic Parameters from 2D Human Pose Estimation
@@
IV-C1 Initial Estimation
Given the video of a human performing actions in front of a mirror, we first apply a state-of-the-art 2D human pose estimator on each frame to get the 2D joint positions of the real human and the mirrored human. The 2D joint locations on each joint are then matched and used in calibration. Special care must be taken, since the joints of the real human and the mirrored human are symmetrically associated. For instance, the right ankle joint of the real human should correspond to the left ankle of the mirrored human. We employ the modified eight-point algorithm in Section IV-A to calculate the reflective fundamental matrix and the reflective essential matrix. Lastly, the reflective essential matrix is used to decompose and calculate the virtual camera pose with respect to the real camera following the procedure in the Appendix.

### v3

@@ Abstract
In this paper, we present the novel task of estimating the extrinsic parameters of a virtual camera relative to a real camera in exercise videos with a mirror. This task poses a significant challenge in scenarios where the views from the real and mirrored cameras have no overlap or share salient features. To address this issue, prior knowledge of a human body and 2D joint locations are utilized to estimate the camera extrinsic parameters when a person is in front of a mirror. We devise a modified eight-point algorithm to obtain an initial estimation from 2D joint locations. The 2D joint locations are then refined subject to human body constraints. Finally, a RANSAC algorithm is employed to remove outliers by comparing their epipolar distances to a predetermined threshold. MirrorCalib achieves a rotation error of 1.82° and a translation error of 69.51 mm on a collected real-world dataset, which outperforms the state-of-art method.
1 Introduction
@@
Estimating 3D human motion in exercise videos is crucial to many applications including but not limited to the assessment of trainee exercise quality [36], virtual or augmented reality [37], and motion captioning [38, 39]. However, reconstructing 3D human motion from monocular 2D video input is an ill-posed problem, due to depth ambiguity. The inclusion of mirror view in exercise videos provides an additional perspective for the instructors. If the geometric relationship between the real view and the virtual view can be established, in principle, it becomes feasible to determine 3D joint positions through triangulation.
@@
Compared to existing work that calibrates virtual cameras [6, 7, 8, 9], our work differs in two key aspects. First, previous studies [6, 7, 8, 9] often calibrate virtual cameras in catadioptric stereo systems using multiple mirrors. Such systems are not common in daily use. In contrast, MirrorCalib only requires the view of one fixed mirror, which is widely available in gyms and dance studios. Second, since earlier work considers scenarios with more than one mirror, it is easier to obtain overlapping views and rely on visible points to establish correspondences. In contrast, our work utilizes (generally invisible) 2D human joint locations as feature correspondences. Furthermore, although existing work calibrates the extrinsic parameters of real cameras using 2D human joints [11, 12], after initial estimation, extrinsic parameters and 3D human joints are optimized as separate variables. MirrorCalib, in contrast, jointly optimizes the two through the inclusion of epipolar constraints on human joints.
The rest of the paper is organized as follows: a presentation of the necessary background for our approach is provided in Section 2. In Section 3, a detailed explanation of MirrorCalib is provided. Experiment setup, datasets and evaluation results are presented in Section 4 followed by conclusion in Section 5.
@@
2 Background
@@
In this section, key geometric concepts relevant to estimating virtual camera poses with respect to a real camera are provided.
@@
2.1 Virtual Camera Projection with Reflective Surfaces and Reflective Epipolar Geometry
In this paper, we treat the real camera frame as the coordinate system where the optical center of the real camera is the origin. Let the intrinsic parameter matrix of the camera be $\mathbf{K}$. We consider the configuration depicted in Figure 2, in which a perspective camera $C$ is in front of a planar mirror $\pi$ that is uniquely defined by its normal vector $\mathbf{n}$ and a distance $d$ to $C$.
As shown in Figure 2, the projection $\mathbf{u^{\prime}}$ of the mirrored point $p^{\prime}$ can be seen as the projection $\mathbf{u}^{\prime\prime}$ of the point $p$ on the image plane of a virtual camera $C^{\prime}$. The reference frame of the virtual camera is obtained by reflecting that of the real camera about the mirror $\pi$. It should be noted that the real camera is right-handed i.e., the orientation of the camera’s coordinate system follows the right-hand rule, while the virtual camera is left-handed. Since both the real camera and its virtual counterpart share the same set of intrinsic parameters $\mathbf{K}$, the purpose of camera calibration in this context is to estimate the extrinsic parameters between these two camera frames, specifically, the rotation matrix $\mathbf{R}$ and the translation vector $\mathbf{t}$.
Denote by $\tilde{\mathbf{u}}$ and $\tilde{\mathbf{u}}^{\prime}$ the homogeneous coordinates of the point $p$ and the point $p^{\prime}$ projected onto the image planes of the real camera $C$. From [16], $\tilde{\mathbf{u}}$ and $\tilde{\mathbf{u}}^{\prime}$ satisfy the reflective epipolar constraint:
$(\tilde{\mathbf{u}}^{\prime})^{\top}\mathbf{F}\tilde{\mathbf{u}}=0,$ (1)
@@ 2.1 Virtual Camera Projection with Reflective Surfaces and Reflective Epipolar Geometry
$\mathbf{F}\triangleq\mathbf{K}^{\top}\mathbf{E}\mathbf{K}^{-1}$ (2)
@@
is the reflective essential matrix, and $[\mathbf{n}]_{\times}$ is the skew-symmetric matrix associated with the normal vector $\mathbf{n}$. Note that since the reflective essential matrix is a skew-symmetric matrix, after transformation by the intrinsic camera matrix, the reflective fundamental matrix is also a skew-symmetric matrix.
3 Methodology
@@
3.1 Eight-Point Algorithm for Solving Reflective Fundamental Matrix
Typically a pair of real cameras have nine unknowns in the fundamental matrix. However, in the case of the reflective fundamental matrix between a real camera and its virtual one, as explained in Section 2.1, the fundamental matrix is a skew-symmetric matrix determined by only three unknowns. Therefore, the transformed fundamental matrix after image normalization [17] is:
@@ 3.1 Eight-Point Algorithm for Solving Reflective Fundamental Matrix
where $[x_{1},x_{2}...x_{6}]$ are unknowns. The normalized fundamental matrix can be solved through least-square minimization. Specifically, the matrix $\mathbf{\hat{A}}$ is obtained by forming a row vector for each pair of matching points $[-\hat{v}^{\prime}\hat{u}+\hat{u}^{\prime}\hat{v},\hat{u}^{\prime},\hat{v}^{\prime},\hat{u},\hat{v},1]$, where $\hat{u}$’s and $\hat{v}$’s are normalized coordinates. The objective is to minimize:
@@
where $\mathbf{f’}$ ’ is the vector $[x_{1},x_{2}...x_{6}]^{\intercal}$ that determines $\mathbf{F}^{\prime}$. The least-square solution to Equation (5) can be calculated by getting the smallest eigenvalue of $\mathbf{\hat{A}^{\top}\hat{A}}$. After $\mathbf{F^{\prime}}$ is obtained, we transform it to the original reflective fundamental matrix $\mathbf{F}$ by the following Equation (6), where $\mathbf{T^{\prime}}$ and $\mathbf{T}$ are normalization matrix of real view and its virtual view.
$\mathbf{F}=\mathbf{T^{\prime}}^{\top}\mathbf{F^{\prime}}\mathbf{T},$ (6)
3.2 Reflective Essential Matrix Decomposition
@@
Similar to a conventional essential matrix, $\mathbf{R}$ and $\mathbf{t}$ can be obtained from the reflective essential matrix $\mathbf{E}$ using singular value decomposition (SVD).
3.3 Estimating Virtual Camera Extrinsic Parameters from 2D Human Pose Estimation
@@
3.3.1 Initial Estimation
Given the video of a human performing actions in front of a mirror, we first apply a state-of-the-art 2D human pose estimator on each frame to get the 2D joint positions of the real human and the mirrored human. The 2D joint locations on each joint are then matched and used in calibration. Special care must be taken, since the joints of the real human and the mirrored human are symmetrically associated. For instance, the right ankle joint of the real human should correspond to the left ankle of the mirrored human. We employ the modified eight-point algorithm in Section 3.1 to calculate the reflective fundamental matrix and the reflective essential matrix. Lastly, the reflective essential matrix is used to decompose and calculate the virtual camera pose with respect to the real camera using singular value decomposition (SVD).